# Quantifying the robustness of the Auriga galaxy formation model

### v1

@@ 6 Intrinsic variance and numerical resolution
Having quantified the intrinsic variance of the Auriga model at fixed numerical resolution, an obvious next step is to extend this comparison to simulations with different numerical resolution. In Figure 10 we compare two sets of simulations with $7$ realisations each. We use the realisations discussed in this paper at the standard resolution of Auriga (L4), and a set of simulations with $8$x worse mass resolution (L5) for all mass components, that is for gas, stars, and dark matter. We compare star formation histories (left panels of Figure 10) and profiles of the stellar mass surface density at $z=0$ (right panels of Figure 10). The bands show interpolated $16\%$ and $84\%$ percentiles. We immediately see that for the first $10\,\mathrm{Gyr}$ of cosmic evolution the star formation rate is systematically lower by $\approx 20\%$ in the simulations with worse mass resolution than in the standard resolution simulations. The difference is much larger than the intrinsic variability at fixed resolution. This difference is consistent with previous results for the Auriga model (Grand et al., 2017, 2021) and similarly the IllustrisTNG model (Pillepich et al., 2018). In the last few Gyr, when the star formation rate drops significantly in all simulations and the relative fractional difference between simulations at the same resolution increases (see also Figure 4), the distributions overlap. However, the mean is still shifted systematically to lower star formation rates at lower resolution.
@@ 7 Summary and discussion
Finally, we showed in Figure 10 and Figure 11 that changing the numerical resolution of the simulation leads to significant systematic changes, which are much bigger than the intrinsic variability of the Auriga model at fixed resolution. The small intrinsic variance of the Auriga model makes it attractive to study physical model variations at fixed resolution (see, e.g. Pakmor et al., 2017; Buck et al., 2020; Rodríguez Montero et al., 2024; van de Voort et al., 2021). However, the systematic changes to the galaxy properties with changing numerical resolution not only limits proper resolution studies, but also limits going to higher resolution without recalibrating a model.
Recent studies looking into the intrinsic variability of galaxy formation models have shown that it is important to understand the robustness of a model to interpret its results, and that the robustness is significantly different for different models (Keller et al., 2019; Davies et al., 2021; Borrow et al., 2023). We suspect that the somewhat smaller variability of the Auriga galaxy formation model compared to these other models is a result of the relatively low importance of individual random numbers. On the opposite end, the blastwave feedback (Keller et al., 2019) in which local feedback immediately changes the local environment around the event because it disables cooling there, seems to have the largest variation.
This should motivate us to include similar robustness and convergence studies when designing new galaxy formation models, in particular for cosmological zoom simulations. Ideally this would allow us to increase the robustness of a model not only at fixed resolution, but also when changing the numerical resolution. As one specific consequence of the study presented here, we will introduce a new approach that improves the stellar mass resolution in cosmological galaxy simulations, but avoids systematically changing galaxy properties at the same time (Pakmor et al. in prep).

### v2

@@ 6 Intrinsic variance and numerical resolution
Having quantified the intrinsic variance of the Auriga model at fixed numerical resolution, an obvious next step is to extend this comparison to simulations with different numerical resolution. In Figure 11 we compare two sets of simulations with $7$ realisations each. We use the realisations discussed in this paper at the standard resolution of Auriga (L4), and a set of simulations with 88x worse mass resolution (L5) for all mass components, that is for gas, stars, and dark matter. We compare star formation histories (left panels of Figure 11) and profiles of the stellar mass surface density at $z=0$ (right panels of Figure 11). The bands show interpolated $16\%$ and $84\%$ percentiles. We immediately see that for the first $10\,\mathrm{Gyr}$ of cosmic evolution the star formation rate is systematically lower by $\approx 20\%$ in the simulations with worse mass resolution than in the standard resolution simulations. The difference is much larger than the intrinsic variability at fixed resolution. This difference is consistent with previous results for the Auriga model (Grand et al., 2017, 2021) and similarly the IllustrisTNG model (Pillepich et al., 2018). In the last few Gyr, when the star formation rate drops significantly in all simulations and the relative fractional difference between simulations at the same resolution increases (see also Figure 4), the distributions overlap. However, the mean is still shifted systematically to lower star formation rates at lower resolution.
@@
The IllustrisTNG model also shows a reduction of variability at better numerical resolution when comparing pairs of Milky Way-like galaxies in different realisations (Genel et al., 2019). However, this reduction of variability mostly disappears for lower mass galaxies. Moreover, Genel et al. (2019) demonstrate that for sufficiently high numerical resolution the variability in the IllustrisTNG model for these low-mass galaxies is not primarily driven by random numbers in the subgrid model.
@@ 7 Summary and discussion
Finally, we showed in Figure 11 and Figure 12 that changing the numerical resolution of the simulation by a factor of $8$ leads to significant systematic changes, which are much bigger than the intrinsic variability of the Auriga model at fixed resolution. The small intrinsic variance of the Auriga model makes it attractive to study physical model variations at fixed resolution (see, e.g. Pakmor et al., 2017; Buck et al., 2020; Rodríguez Montero et al., 2024; van de Voort et al., 2021). However, the systematic changes to the galaxy properties with changing numerical resolution not only limits proper resolution studies, but also limits going to higher resolution without recalibrating a model.
Recent studies looking into the intrinsic variability of galaxy formation models have shown that it is important to understand the robustness of a model to interpret its results, and that the robustness is significantly different for different models (Keller et al., 2019; Genel et al., 2019; Davies et al., 2021; Borrow et al., 2023). A detailed quantitative comparison between the intrinsic variability of different models is hard because most published work limits quantitative discussions to the total stellar mass.
The four different feedback models discussed in Keller et al. (2019) show significant differences. They focus on the evolution of stellar mass and show that their "Superbubble" feedback has the largest variance in stellar mass of $10\%$ at $z=2$ over $128$ realisations and of $20\%$ between a pair of simulations evolved further to $z=0.8$. For this model they also find differences in the stellar surface density at $z=0.8$ of almost an order of magnitude. This level of variance is significantly larger than for the Auriga model. Their other feedback models have comparable variance at $z=2$ and lower variance at $z=0.8$, but at the cost of inefficient feedback that is essentially consistent with no feedback.
The comparison of a set of large volume simulations with the IllustrisTNG model (Genel et al., 2019) showed that massive galaxies can end up with completely different morphologies. This is a result of the very stochastic radio-mode AGN feedback model of IllustrisTNG. The Auriga model, which is very similar to the IllustrisTNG model in terms of stellar feedback, features a radio-mode AGN feedback model that is much more gentle and less stochastic (Grand et al., 2017). The IllustrisTNG simulations also show roughly $5\%$ variance in the value of the maximum of the circular velocity curve, averaged over about $100$ Milky Way-like galaxies. This is roughly consistent with our results for the Auriga model. The stellar mass of matched Milky Way-like galaxies between the IllustrisTNG boxes varies by $10\%$ at $z=0$ for the highest resolution box with roughly comparable mass resolution to the standard Auriga (L4) resolution, slightly larger than the variance ($\lesssim 5\%$) we find for our single galaxy.
Similar studies with the EAGLE model (Davies et al., 2021) and the SWIFT-EAGLE model (Borrow et al., 2023) show variance on the $1\%$ level in the halo mass at $z=0$ for Milky Way-like galaxies, similar to our simulations. Moreover, they show an average variance in the stellar mass of Milky Way-like galaxies on the $10\%$ level (EAGLE) and between $5\%$ and $10\%$ (SWIFT-EAGLE), again slightly larger than for our single galaxy.
These differences should motivate us to include similar robustness and convergence studies when designing new galaxy formation models, in particular for cosmological zoom simulations. Ideally this would allow us to increase the robustness of a model not only at fixed resolution, but also when changing the numerical resolution. As one specific consequence of the study presented here, we will introduce a new approach that improves the stellar mass resolution in cosmological galaxy simulations, but avoids systematically changing galaxy properties at the same time (Pakmor et al., 2025).